# A Survey of AI Scientists: Surveying the automatic Scientists and Research

### v1

@@ Abstract
A new scientific paradigm, the AI Scientist, has coalesced at the intersection of artificial intelligence and epistemology, promising a fundamental shift from AI-assisted analysis to end-to-end autonomous discovery. Catalyzed by rapid advances in large language models, multi-agent orchestration, and robotic automation, these systems are architected to emulate the complete scientific workflow—from initial hypothesis generation to the final synthesis of publishable findings. This transition moves beyond using AI as an instrument of inquiry, positioning it as a potential originator of scientific knowledge. However, the rapid and unstructured proliferation of these systems has created a fragmented research landscape, obscuring overarching methodological principles and developmental trends. This survey provides a systematic and comprehensive synthesis of this emerging domain by introducing a unified, six-stage methodological framework that deconstructs the scientific process into: Literature Review, Idea Generation, Experimental Preparation, Experimental Execution, Scientific Writing, and Paper Generation. Through this analytical lens, we systematically map and analyze dozens of seminal works from 2022 to late 2025, revealing a clear three-phase evolutionary trajectory. Our analysis charts the field’s progression from an initial phase of Foundational Modules (2022–2023), focused on task-specific automation, through a period of Closed-Loop Integration (2024), to the current frontier of Scalability, Impact, and Collaboration (2025–present). By synthesizing these developments, this survey identifies key architectural patterns and highlights the dual research thrusts toward both greater machine autonomy and more sophisticated human-in-the-loop synergy. We conclude by presenting a forward-looking agenda that addresses critical open challenges in robustness, generalizability, and ethical governance. Ultimately, this work provides a critical roadmap for the field, intended to guide the next generation of systems toward becoming trustworthy, verifiable, and indispensable partners in human scientific inquiry. Project Github: https://github.com/Mr-Tieguigui/Survey-for-AI-Scientist.
@@ 1 Introduction
This survey focuses on research between 2022 and 2025 that enables autonomous or near-autonomous execution of the scientific method. To ensure conceptual clarity and structural coherence, we adopt the radial organization illustrated in Figure 1, which conceptualizes the AI Scientist as a core engine driving a cyclical workflow that is instantiated across various scientific domains. At its heart lies the central concept of the Automatic AI Scientist System. Surrounding this core is an inner ring depicting the six fundamental methodological categories that constitute the scientific workflow: (1) Literature Review focuses on extracting, structuring, and reasoning over scientific corpora to establish prior knowledge foundations [27]; (2) Idea Generation addresses hypothesis formation and creative scientific reasoning [25, 26]; (3) Experimental Preparation encompasses data selection, variable definition, simulation setup, and statistical initialization [29, 30]; (4) Experimental Execution integrates protocol orchestration, robotic control, feedback-based iteration, and tool-use reasoning [12, 11, 31]; (5) Scientific Writing involves multi-modal evidence organization, figure and table generation, and result articulation [27, 32]; and (6) Paper Generation targets the synthesis of publishable, verifiable manuscripts that combine reasoning, visualization, and scholarly consistency [21, 33]. As shown in Figure 1, encircling this workflow is an outer ring representing the application layers, showcasing where these methodologies are applied—from general-purpose frameworks to specialized domains such as Chemistry, Biology, Physics, and Meta-Science. This concentric structure provides a unified, navigable map for understanding how a core autonomous system executes a fundamental scientific process across a diverse and expanding landscape of applications.
Figure 1: The concentric architecture of an AI Scientist System. The core represents the central Automatic AI Scientist System engine. This engine drives the six-stage methodological workflow, depicted in the inner ring, which forms a continuous, iterative cycle from Literature Review to Paper Generation. The outer ring illustrates the key application domains where this workflow is instantiated, from general-purpose frameworks to specialized fields like Chemistry, Biology, and Physics.

### v2

@@ Abstract
A new scientific paradigm, the AI Scientist, has coalesced at the intersection of artificial intelligence and epistemology, promising a fundamental shift from AI-assisted analysis to end-to-end autonomous discovery. Catalyzed by rapid advances in large language models, multi-agent orchestration, and robotic automation, these systems are architected to emulate the complete scientific workflow—from initial hypothesis generation to the final synthesis of publishable findings. This transition moves beyond using AI as an instrument of inquiry, positioning it as a potential originator of scientific knowledge. However, the rapid and unstructured proliferation of these systems has created a fragmented research landscape, obscuring overarching methodological principles and developmental trends. This survey provides a systematic and comprehensive synthesis of this emerging domain by introducing a unified, six-stage methodological framework that deconstructs the scientific process into: Literature Review, Idea Generation, Experimental Preparation, Experimental Execution, Scientific Writing, and Paper Generation. Through this analytical lens, we systematically map and analyze dozens of seminal works from 2022 to late 2025, revealing a clear three-phase evolutionary trajectory. Our analysis charts the field’s progression from an initial phase of Foundational Modules, focused on task-specific automation, through a period of Closed-Loop Integration, to the current frontier of Scalability, Impact, and Collaboration. By synthesizing these developments, this survey identifies key architectural patterns and highlights the dual research thrusts toward both greater machine autonomy and more sophisticated human-in-the-loop synergy. We conclude by presenting a forward-looking agenda that addresses critical open challenges in robustness, generalizability, and ethical governance. Ultimately, this work provides a critical roadmap for the field, intended to guide the next generation of systems toward becoming trustworthy, verifiable, and indispensable partners in human scientific inquiry. Project Github: https://github.com/Mr-Tieguigui/Survey-for-AI-Scientist.
@@ 1 Introduction
This survey focuses on research between 2022 and 2025 that enables autonomous or near-autonomous execution of the scientific method. To ensure conceptual clarity and structural coherence, we adopt the comprehensive architectural landscape illustrated in Figure 1 as the central organizing principle for our analysis. This structure is defined by two primary axes. The horizontal axis delineates the six sequential methodological stages of the scientific workflow: (1) Literature Review focuses on extracting, structuring, and reasoning over scientific corpora to establish prior knowledge foundations [27]; (2) Idea Generation addresses hypothesis formation and creative scientific reasoning [25, 26]; (3) Experimental Preparation encompasses data selection, variable definition, simulation setup, and statistical initialization [29, 30]; (4) Experimental Execution integrates protocol orchestration, robotic control, feedback-based iteration, and tool-use reasoning [12, 11, 31]; (5) Scientific Writing involves multi-modal evidence organization, figure and table generation, and result articulation [27, 32]; and (6) Paper Generation targets the synthesis of publishable, verifiable manuscripts that combine reasoning, visualization, and scholarly consistency [21, 33]. The vertical axis defines four distinct layers of abstraction in a top-down hierarchy, flowing from high-level Application Domains, through concrete Scientific Tasks & Products and enabling System Architectures & Methods, down to the foundational Models & Capabilities. This 4x6 matrix provides a unified map for situating any given work by cross-referencing its methodological focus with its layer of contribution. Complementing this matrix, the figure also includes a timeline at the bottom that summarizes the field’s three-phase historical evolution, providing a chronological perspective on the progression from foundational modules to fully integrated systems.
Figure 1: The Architectural Landscape of the Survey of AI Scientist for Automatic Research. The main 4x6 matrix maps the six methodological stages of the scientific workflow (horizontal axis) against four top-down layers of abstraction, from Applications to Models (vertical axis). The panel at the bottom illustrates the field’s three-phase historical evolution, categorizing representative works to provide a chronological perspective on the development of AI Scientist systems.